OpenStudy (anonymous):

What is the product of the polynomials. (5x^2-2x+4) (2x^2+3x+1)

OpenStudy (anonymous):

I got 10x^4-6x^2-4. Is this correct?

OpenStudy (mathstudent55):

No. The x^3 term and x term are missing, and some therms are incorrect.

OpenStudy (anonymous):

How do I get this term?

OpenStudy (mathstudent55):

Multiply every term of the first polynomial by every term of the second polynomial. After that step, you should have 9 terms. Then collect like terms.

OpenStudy (anonymous):

10x^4+15x^3+5x^2

OpenStudy (anonymous):

This is confusing -_-

OpenStudy (mathstudent55):

\( (5x^2-2x+4)(2x^2+3x+1) \) \(= ~~5x^2 \times 2x^2 + 5x^2 \times 3x + 5x^2 \times 1 \) \(~~ - 2x \times 2x^2 - 2x \times 3x - 2x \times 1 \) \(~~ + 4 \times 2x^2 + 4 \times 3x + 4 \times 1\)

OpenStudy (mathstudent55):

Now you do all those multiplications, then you collect like terms.

OpenStudy (anonymous):

30^6-12^2+24^2

OpenStudy (anonymous):

Is that correct @mathstudent55

OpenStudy (mathstudent55):

No. Here is the next step after all the multiplications are done. 10x^4 + 15x^3 + 5x^2 - 4x^3 - 6x^2 - 2x + 8x^2 + 12x + 4

OpenStudy (mathstudent55):

Now you need to collect like terms.

OpenStudy (anonymous):

22^4+11^3+7x^2.

OpenStudy (mathstudent55):

I only see one term with x^4; it's 10x^4. How do you get 22x^4?

OpenStudy (anonymous):

I added all the x^4 up.

OpenStudy (mathstudent55):

Like terms are terms that have the same variable part. That means the same variable and the same exponent.

OpenStudy (mathstudent55):

How can you add all the x^4 terms up if there is only one?

OpenStudy (anonymous):

About 12x^4?

OpenStudy (mathstudent55):

10x^4 + 15x^3 + 5x^2 - 4x^3 - 6x^2 - 2x + 8x^2 + 12x + 4 ^^^^ only x^4 term is this one, 10x^4

OpenStudy (anonymous):

Oh!

OpenStudy (mathstudent55):

That means you start collecting like terms by writing the x^4 term: 10x^4 Now move on to the next lower term, x^3

OpenStudy (anonymous):

10x^4+11^3+7x^2+4?

OpenStudy (anonymous):

Im confused on the 12x+4.

OpenStudy (mathstudent55):

10x^4 + 15x^3 + 5x^2 - 4x^3 - 6x^2 - 2x + 8x^2 + 12x + 4 ^^^^^ ^^^^ x^3 terms are + 15x^3 and - 4x^3. Collect them together and get: 11x^3, so you write that next: 10x^4 + 11x^3

OpenStudy (mathstudent55):

We'll get there.

OpenStudy (anonymous):

10x^4+11x^3+7x^2.

OpenStudy (mathstudent55):

Now we move on to the next lower power of x, x^2. Let's look for all x^2 terms: 10x^4 + 15x^3 + 5x^2 - 4x^3 - 6x^2 - 2x + 8x^2 + 12x + 4 ^^^^ ^^^^^ ^^^^^ We collect together 5x^2 - 6x^2 + 8x^2 and we get +7x^2, so we write it next: 10x^4 + 11x^3 + 7x^2

OpenStudy (mathstudent55):

Now that the x^2 terms are done, we move on to the next lower power of x, x to the 1 power, or just plain x. 10x^4 + 15x^3 + 5x^2 - 4x^3 - 6x^2 - 2x + 8x^2 + 12x + 4 ^^^ ^^^^ The x terms are -2x + 12x. We add them together and get: 10x. We write that next: 10x^4 + 11x^3 + 7x^2 + 10x

OpenStudy (anonymous):

About the 4?

OpenStudy (mathstudent55):

The last step is the term with no x, called the constant term. There is only one of them, it's + 4, so we write that last: 10x^4 + 11x^3 + 7x^2 + 10x + 4

OpenStudy (mathstudent55):

That is the final answer.

OpenStudy (anonymous):

Thank you so much.

OpenStudy (mathstudent55):

This is what you need to do for all these problems of multiplying polynomials together.